# Advances in Feed-Forward 3D Reconstruction and View Synthesis: A Survey

### v1

@@ 2.1 NeRF
Neural radiance fields (NeRF) [17] has recently gained significant attention for high-quality novel view synthesis using implicit scene representations and differentiable volume rendering. By leveraging MLPs, NeRF reconstructs 3D scenes from multiview 2D images, enabling the generation of novel views with excellent multiview consistency. However, a major limitation of NeRF is its requirement for per-scene optimization, which restricts its generalization to unseen scenes. To address this, feed-forward approaches have been proposed, where neural networks learn to infer NeRF representations directly from sparse input views, thereby eliminating the need for scene-specific optimization. As a pioneering feed-forward NeRF work, PixelNeRF [22] introduces a conditional NeRF framework that leverages pixel-aligned image features extracted from input images, allowing the model to generalize across diverse scenes and perform novel view synthesis from sparse observations. A large number of follow-ups adopt various techniques for feed-forward NeRF, and we broadly categorize them into the following categories based on feature representations.
@@ 2.2 Pointmap
To improve computational efficiency, a couple of workers introduce a memory mechanism that aims to incrementally process the input and add points to a canonical 3D space by incrementally updating a scene’s latent state. Spann3R [62] introduces a spatial memory network, enabling multiview input and improving efficiency to eliminate the need for global alignment. MUSt3R [63] extends the DUSt3R architecture by introducing a symmetric design and a memory mechanism, effectively reducing computational complexity when handling multiview inputs. Closely, CUT3R [3] proposes a Continuous Updating Transformer that simultaneously updates the state with new information and retrieves the information stored in the state. This formulation is general and able to handle both video and photo collections, and process both static and dynamic scenes. However, with the increased number of processed frames, memory-based methods face capacity constraints, which can result in the degradation or loss of information from earlier frames. To address this issue, Point3R [64] takes inspiration from the human memory mechanism and proposes a spatial pointer memory, where each pointer is anchored at a 3D position and links to a dynamically evolving spatial feature. Besides, Driv3R [65] extends the memory mechanism to support efficient temporal integration, enabling large-scale dynamic scene reconstruction from sequences of multiview inputs.
@@
As a promising and powerful foundation for 3D reconstruction, VGGT [2] presents a large feed-forward transformer-based architecture that directly predicts all essential 3D attributes, such as camera intrinsics and extrinsics, point maps, depth maps, and 3D point tracks, without the need for post-processing, leading to state-of-the-art 3D point and camera pose reconstruction.
@@ 2.3.1 Gaussian Image
To improve the reconstruction quality, several methods are subsequently proposed to leverage large models with strong capacity to learn generic scene priors from large-scale datasets for 3D scene reconstruction. Based on the 3D large reconstruction model (LRM) [10] that achieves impressive sparse-view 3D object reconstruction by learning general reconstruction priors from extensive datasets of 3D objects, GRM [78] directly maps input image pixels to a set of pixel-aligned 3D Gaussians for feed-forward 3DGS-based object reconstruction. Flash3D [79] introduces a high-quality depth predictor as prior to achieve single-view scene-level reconstruction. Concurrently, GS-LRM [80] incorporates Transformer-based LRM to formulate per-pixel Gaussian prediction as a sequence-to-sequence mapping and achieves remarkable performance across both objects and scenes. This research line is further extended by eFreeSplat [81] that leverages a large vision transformer encoder [52] as 3D priors for Gaussian image prediction, Long-LRM [82] that combines Mamba2 blocks [83] with Transformer layers to handle long sequence of input images, and FreeSplatter [84] that jointly predicts Gaussian images and estimates camera poses. On the other hand, these works are constrained to reconstruct existing image observations without generative ability. To address it, LGM [85] introduces pre-trained diffusion models [86, 12, 87, 88] to generate multiview images, followed by large multiview Gaussian models to predict multiview Gaussian images. Wonderland [89] further leverages a pre-trained video diffusion model [90] to generate informative video latents from a single image for pixel-aligned 3DGS prediction.
@@ 2.3.2 Gaussian Volume
Gaussian volume [74] represents 3D with Gaussian voxel grids, where each voxel comprises multiple Gaussian primitives. A typical feed-forward 3DGS method using a Gaussian volume representation is LaRa [74], which aims to reduce the heavy training cost associated with 360° bounded radiance field reconstruction. As shown in Fig. 5 (b), it first builds 3D features and embedding volumes and then leverages a volume transformer to reconstruct a Gaussian volume, enabling progressively and implicitly feature matching and leading to higher quality results and faster convergence. GaussianCube [102] proposes a structured and explicit radiance representation for 3D object generation from a single image. Besides, to enable Gaussian densification in feed-forward 3DGS, GD [103] builds upon LaRa and introduces a generative densification that exploits prior knowledge from large multiview datasets and densifies feature representations from feed-forward 3DGS. SCube [104] further advances large-scale scene reconstruction by introducing VoxSplat, a high-resolution sparse-voxel Gaussian representation generated via a hierarchical latent diffusion model conditioned on sparse posed images.
@@ 2.4.1 Mesh
To generate artist-created meshes with high-quality topology, several methods [123, 124, 125] draw inspiration from large language models, treating 3D meshes as sequences and introducing autoregressive transformer architectures tailored to this sequential representation. Specifically, MeshGPT [123] leverages VQVAE [126] to learn a mesh vocabulary and employs a decoder-only transformer to autoregressively generate triangle meshes in a sequential manner. To mitigate cumulative errors inherent in VQVAE-based sequence representations, MeshXL [124] introduces a neural coordinate field for sequential 3D mesh representation, enabling high-quality autoregressive mesh generation. However, these methods struggle to learn the shape distribution and the topology distribution simultaneously. To address this issue, MeshAnything [125] introduces shape-conditioned artist-created mesh generation. It leverages a pre-trained encoder [127] to extract shape features, which are then injected into the VQVAE-based sequence representation, eliminating the need to learn shape distribution and enabling the model to focus solely on topology distribution learning.
@@ 2.5.2 Generative Feed-Forward View Synthesis
Earlier works primarily use transformer-based autoregressive models [126, 148]. For example, GFVS [149] approaches novel view synthesis from a single view by treating it as sampling target images from a learned distribution conditioned on a source image and camera transformation, where the distribution is modeled autoregressively by leveraging a VQGAN [148] with a transformer. ViewFormer [150] extends single-view NVS of GFVS to multiview NVS. Specifically, it first uses a VQVAE codebook [126] to encode images into latent representations, then queries latent codes of target views and employs a transformer to map the latent codes to image tokens, which are subsequently decoded into novel views.
@@
Image Diffusion Model. Zero-1-to-3 [117] leverages the latent diffusion model [92] pretrained for text-to-image generation and replaces text embedding with relative camera poses as conditioning to achieve novel view synthesis as illustrated in Fig. 7(a). ZeroNVS [151] extends Zero-1-to-3 to achieve single-view scene-level novel view synthesis by finetuning it on diverse large-scale object and real-scene datasets [27, 14, 152]. However, these methods still face challenges in generating consistent novel views. To address this limitation, SyncDreamer [153] initializes the diffusion model with pretrained Zero-1-to-3 weights and extends the diffusion model to capture the joint probability distribution of multiview images, generating novel views with multiview consistency. Zero123++ [12] arranges six surrounding views into a single image, facilitating accurate joint distribution modeling of an object’s multiview representations. Consistent123 [154] combines Zero-1-to-3 and stable diffusion to provide diffusion priors to ensure the multiview-consistent synthesized novel views. ConsistNet [155] builds on Zero-1-to-3 as its backbone and performs multiple diffusions in parallel, each handling a specific viewpoint. To enforce multiview geometric consistency, it incorporates a dedicated plug-in block that aligns the generated images accordingly. MVDream [87] proposes a multiview diffusion model that leverages both 2D and 3D data, combining the generalizability of 2D diffusion models with the consistency of 3D renderings. It further demonstrates that a multiview diffusion model can implicitly provide generalizable 3D priors without specific 3D representations.
Video Diffusion Model. Video diffusion models [156, 157] have achieved impressive realism in video synthesis and are thought to inherently and implicitly capture 3D structures. Building on this capability, recent approaches have explored leveraging their priors to generate multiview images for high-quality 3D reconstruction. For example, ReconX [147], as illustrated in Fig. 7(b), harnesses the powerful generative prior of large pretrained video diffusion models [157] to synthesize novel views. It encodes extracted point clouds as 3D structural conditions, ensuring multiview consistency in the generated novel views. Similarly, ViewCrafter [8] first builds a point cloud representation and then uses point cloud renderings as the conditions of the video diffusion model [157] to enable consistent and accurate novel view synthesis. MultiDiff [158] leverages a single reference image and a predefined target camera trajectory as conditions, utilizing depth cues to encourage consistent novel view synthesis. More recently, GenFusion [159], DifFusion3D+ [160], and SpatialCrafter [161] have bridged reconstruction and generation by using video diffusion models as scene reconstruction refiners, enabling both artifact removal and scene content expansion.
@@ 3 Tasks & Applications
Several earlier methods employ voxel-based representations (e.g., PlatonicGAN [162]) or 3D feature representations (e.g., HoloGAN [163] and BlockGAN [164]). However, these approaches often suffer from limited multiview consistency. To address this, GRAF [165] introduces a generative radiance field as a 3D representation, significantly improving consistency across different viewpoints. It designs a conditional NeRF that takes shape and appearance latent codes as conditions and produces images via volume rendering. PiGAN [166] leverages implicit neural representations with periodic activation functions to model scenes as view-consistent radiance fields. Subsequently, GIRAFFE [167] constructs compositional generative radiance fields for scene representations, enabling controllable image synthesis. StyleNeRF [168] combines NeRF-based 3D representations with a style-based generative model for high-resolution, 3D-consistent image synthesis. EG3D [43] introduces an explicit-implicit triplane representation to achieve efficient and high-quality 3D-aware image synthesis.
@@ 3.3 Pose-free 3D Reconstruction
The development of feed-forward models has enabled the reconstruction of 3D scenes from unposed images or videos without the need for per-scene optimization. FlowCam [187] employs a single-view feed-forward generalizable NeRF to generate point maps for different input viewpoints, using optical flow to estimate poses and integrate point maps from multiple views to reconstruct neural radiance fields. CoPoNeRF [188] extracts multi-level features from image pairs to build 4D correlation maps capturing pixel-pair similarities. These maps are further refined for flow and pose estimation, enabling color and depth rendering from the updated features and poses.
@@
Additionally, several methods build upon DUSt3R [1] for pose-free 3D reconstruction. DUSt3R itself, as a pioneering feed-forward method, utilizes a transformer-based architecture to regress 3D point maps directly from image pairs. Spann3R [62] augments DUSt3R with a spatial memory network, allowing multiview inputs and improving efficiency by eliminating global alignment. However, Spann3R’s sequential processing introduces error accumulation in reconstruction. Fast3R [59] overcomes this limitation by introducing a global fusion transformer, processing multiple views simultaneously and significantly enhancing reconstruction quality. Conversely, CUT3R [3] refines sequential reconstruction by maintaining and incrementally updating a persistent internal state that encodes scene content. Instead of relying on pairwise feature matching with previous views, CUT3R updates its internal state continuously and utilizes it directly to predict the pointmap of the current view.
@@ 3.4 Dynamic 3D Reconstruction
Compared to static scene reconstruction, dynamic scene reconstruction poses significant challenges mainly due to the presence of moving objects, changing viewpoints, and temporal variations in scene geometry. Extending feedforward 3D reconstruction for dynamic scenarios mainly involves robust pose estimation to mitigate moving object interference, together with dynamic area segmentation for updating changing environments.
@@
In addition to pointmap-based dynamic scene reconstruction, several recent methods based on 3D Gaussian Splatting (3DGS) have also been proposed for feed-forward dynamic reconstruction. L4GM [206] proposes the first 4D reconstruction model that produces animated objects from single-view videos using per-frame 3DGS representation. 4D-LRM [207] builds upon a transformer-based large reconstruction model, leveraging data-driven training for dynamic object reconstruction. It draws inspiration from 4D Gaussian Splatting [208] and reconstructs dynamic objects as anisotropic 4D Gaussian clouds. While prior works focus on dynamic object reconstruction, BulletTimer [6] introduces the first feed-forward model for dynamic scene reconstruction. Building on GS-LRM [80], it incorporates a bullet-time embedding into the input frames and aggregates information across all context frames, enabling feed-forward 3D Gaussian Splatting reconstruction at a specific timestamp. In addition, DGS-LRM [209] introduces the first feed-forward prediction of deformable 3D Gaussians from monocular videos with a transformer-based LRM architecture.
@@ 3.6 Image Matching
Recent advances in feed-forward 3D reconstruction have led to significant progress in image matching. One notable example is MASt3R [4], which builds on the DUSt3R [1] to enable efficient and robust image matching in a single forward pass. By augmenting the DUSt3R architecture with a dedicated head for dense local feature extraction, MASt3R introduces a mechanism to improve matching accuracy while maintaining the robustness characteristic of pointmap-based regression.
However, MASt3R is fundamentally limited to processing image pairs with poor scalability for large image collections. To address this issue, MASt3R-SfM [223] proposes to leverage the frozen encoder of MASt3R for image retrieval, enabling it to process large and unconstrained image collections with quasi-linear complexity in a scalable way. Importantly, the robustness of MASt3R’s local reconstructions allows the SfM pipeline to dispense with traditional RANSAC-based filtering. Instead, optimization is performed through successive gradient-based refinement in both 3D space (via a matching loss) and 2D image space (via reprojection loss), thus highlighting the potential of feed-forward paradigms to serve as both matching engines and geometric optimizers.

### v2

@@ 2.1 NeRF
Neural radiance fields (NeRF) [17] have recently gained significant attention for high-quality novel view synthesis using implicit scene representations and differentiable volume rendering. By leveraging MLPs, NeRF reconstructs 3D scenes from multiview 2D images, enabling the generation of novel views with excellent multiview consistency. However, a major limitation of NeRF is its requirement for per-scene optimization, which restricts its generalization to unseen scenes. To address this, feed-forward approaches have been proposed, where neural networks learn to infer NeRF representations directly from sparse input views, thereby eliminating the need for scene-specific optimization. As a pioneering feed-forward NeRF work, PixelNeRF [22] introduces a conditional NeRF framework that leverages pixel-aligned image features extracted from input images, allowing the model to generalize across diverse scenes and perform novel view synthesis from sparse observations. A large number of follow-ups adopt various techniques for feed-forward NeRF, and we broadly categorize them into the following categories based on feature representations.
@@ 2.2 Pointmap
To improve computational efficiency, a couple of works introduce a memory mechanism that aims to incrementally process the input and add points to a canonical 3D space by incrementally updating a scene’s latent state. Spann3R [62] introduces a spatial memory network, enabling multiview input and improving efficiency to eliminate the need for global alignment. MUSt3R [63] extends the DUSt3R architecture by introducing a symmetric design and a memory mechanism, effectively reducing computational complexity when handling multiview inputs. Closely, CUT3R [3] proposes a Continuous Updating Transformer that simultaneously updates the state with new information and retrieves the information stored in the state. This formulation is general and able to handle both video and photo collections, and process both static and dynamic scenes. However, with the increased number of processed frames, memory-based methods face capacity constraints, which can result in the degradation or loss of information from earlier frames. To address this issue, Point3R [64] takes inspiration from the human memory mechanism and proposes a spatial pointer memory, where each pointer is anchored at a 3D position and links to a dynamically evolving spatial feature. Besides, Driv3R [65] extends the memory mechanism to support efficient temporal integration, enabling large-scale dynamic scene reconstruction from sequences of multiview inputs.
@@
As a promising and powerful foundation for 3D reconstruction, VGGT [2] presents a large feed-forward transformer-based architecture that directly predicts all essential 3D attributes, such as camera intrinsics and extrinsics, point maps, depth maps, and 3D point tracks, without the need for post-processing, leading to state-of-the-art 3D point and camera pose reconstruction. VGGT-Long [73] extends VGGT to handle kilometer-scale sequences through a chunk-and-align pipeline, effectively mitigating its memory constraints in long-sequence handling.
@@ 2.3.1 Gaussian Image
To improve the reconstruction quality, several methods are subsequently proposed to leverage large models with strong capacity to learn generic scene priors from large-scale datasets for 3D scene reconstruction. Based on the 3D large reconstruction model (LRM) [10] that achieves impressive sparse-view 3D object reconstruction by learning general reconstruction priors from extensive datasets of 3D objects, GRM [79] directly maps input image pixels to a set of pixel-aligned 3D Gaussians for feed-forward 3DGS-based object reconstruction. Flash3D [80] introduces a high-quality depth predictor as prior to achieve single-view scene-level reconstruction. Concurrently, GS-LRM [81] incorporates Transformer-based LRM to formulate per-pixel Gaussian prediction as a sequence-to-sequence mapping and achieves remarkable performance across both objects and scenes. This research line is further extended by eFreeSplat [82] that leverages a large vision transformer encoder [52] as 3D priors for Gaussian image prediction, Long-LRM [83] that combines Mamba2 blocks [84] with Transformer layers to handle long sequence of input images, and several pose-free methods, including FreeSplatter [85] and AnySplat [86], that jointly predicts Gaussian images and estimates camera poses, where AnySplat leverages a pretrained VGGT [2] with geometry priors as an auxiliary supervision for the joint prediction. On the other hand, these works are constrained to reconstruct existing image observations without generative ability. To address it, LGM [87] introduces pre-trained diffusion models [88, 12, 89, 90] to generate multiview images, followed by large multiview Gaussian models to predict multiview Gaussian images. Wonderland [91] further leverages a pre-trained video diffusion model [92] to generate informative video latents from a single image for pixel-aligned 3DGS prediction.
@@ 2.3.2 Gaussian Volume
Gaussian volume [75] represents 3D with Gaussian voxel grids, where each voxel comprises multiple Gaussian primitives. A typical feed-forward 3DGS method using a Gaussian volume representation is LaRa [75], which aims to reduce the heavy training cost associated with 360° bounded radiance field reconstruction. As shown in Fig. 5 (b), it first builds 3D features and embedding volumes and then leverages a volume transformer to reconstruct a Gaussian volume, enabling progressively and implicitly feature matching and leading to higher quality results and faster convergence. GaussianCube [104] proposes a structured and explicit radiance representation for 3D object generation from a single image. QuickSplat [105], based on G3R [106], employs neural networks to predict voxel-level features, which are subsequently decoded into Gaussian primitives with MLPs for surface reconstruction. Besides, to enable Gaussian densification in feed-forward 3DGS, GD [107] builds upon LaRa and introduces a generative densification that exploits prior knowledge from large multiview datasets and densifies feature representations from feed-forward 3DGS. SCube [108] further advances large-scale scene reconstruction by introducing VoxSplat, a high-resolution sparse-voxel Gaussian representation generated via a hierarchical latent diffusion model conditioned on sparse posed images.
@@ 2.4.1 Mesh
To generate artist-created meshes with high-quality topology, several methods draw inspiration from large language models, treating 3D meshes as sequences and introducing autoregressive transformer architectures tailored to this sequential representation. Specifically, MeshGPT [127] leverages VQVAE [128] to learn a mesh vocabulary and employs a decoder-only transformer to autoregressively generate triangle meshes in a sequential manner. To mitigate cumulative errors inherent in VQVAE-based sequence representations, MeshXL [129] introduces a neural coordinate field for sequential 3D mesh representation, enabling high-quality autoregressive mesh generation. However, these methods struggle to learn the shape distribution and the topology distribution simultaneously. To address this issue, MeshAnything [130] introduces shape-conditioned artist-created mesh generation. It leverages a pre-trained encoder [131] to extract shape features, which are then injected into the VQVAE-based sequence representation, eliminating the need to learn shape distribution and enabling the model to focus solely on topology distribution learning.
@@ 2.5.2 Generative Feed-Forward View Synthesis
Earlier works primarily use transformer-based autoregressive models [128, 152]. For example, GFVS [153] approaches novel view synthesis from a single view by sampling target images from a learned distribution conditioned on a source image and camera transformation, where the distribution is modeled autoregressively by leveraging a VQGAN [152] with a transformer. ViewFormer [154] further extends single-view NVS of GFVS to multiview NVS by VQVAE codebook [128] and transformer.
@@
Image Diffusion Models. Zero-1-to-3 [121] leverages the latent diffusion model [94] pretrained for text-to-image generation and replaces text embedding with relative camera poses as conditioning to achieve novel view synthesis as illustrated in Fig. 7(a). ZeroNVS [155] extends Zero-1-to-3 to achieve single-view scene-level novel view synthesis by finetuning it on diverse large-scale object and real-scene datasets [27, 14, 156]. However, these methods still face challenges in generating consistent novel views. SyncDreamer [157] addresses this by initializing a diffusion model with Zero-1-to-3 weights and modeling the joint distribution of multiview images. Zero123++[12] arranges six views into a single image for improved multiview representation modeling. ViewDiff [158] also exploits diffusion priors, with emphasis on multi-view consistency in real-world data. Consistent123 [159] combines Zero-1-to-3 and stable diffusion to provide case-aware diffusion priors to ensure the multiview-consistency, while ConsistNet [160] performs parallel viewpoint-specific diffusion and aligns the generated images accordingly to enforce multiview geometric consistency. MVDream [89] proposes a multiview diffusion model that leverages both 2D and 3D data, combining the generalizability of 2D diffusion models with the consistency of 3D renderings. It further demonstrates that a multiview diffusion model can implicitly provide generalizable 3D priors without specific 3D representations.
Video Diffusion Models. Video diffusion models [161, 162] have achieved impressive realism in video synthesis and are thought to inherently and implicitly capture 3D structures. Building on this capability, recent approaches have explored leveraging their priors to generate multiview images for high-quality 3D reconstruction. For example, ReconX [151], as illustrated in Fig. 7(b), harnesses the powerful generative prior of large pretrained video diffusion models [162] to synthesize novel views. It encodes extracted point clouds as 3D structural conditions, ensuring multiview consistency in the generated novel views. Similarly, ViewCrafter [8] first builds a point cloud representation and then uses point cloud renderings as the conditions of the video diffusion model [162] to enable consistent and accurate novel view synthesis. MultiDiff [163] leverages a single reference image and a predefined target camera trajectory as conditions, utilizing depth cues to encourage consistent novel view synthesis. More recently, GenFusion [164], DifFusion3D+ [165], and SpatialCrafter [166] have bridged reconstruction and generation by using video diffusion models as scene reconstruction refiners, enabling both artifact removal and scene content expansion.
@@ 3 Tasks & Applications
Several earlier methods employ voxel-based representations (e.g., PlatonicGAN [167]) or 3D feature representations (e.g., HoloGAN [168] and BlockGAN [169]). However, these approaches often suffer from limited multiview consistency. To address this, GRAF [170] introduces a generative radiance field as a 3D representation, significantly improving consistency across different viewpoints. PiGAN [171] leverages implicit neural representations with periodic activation functions to model scenes as view-consistent radiance fields. Subsequently, GIRAFFE [172] constructs compositional generative radiance fields for scene representations, enabling controllable image synthesis. StyleNeRF [173] combines NeRF-based 3D representations with a style-based generative model for high-resolution, 3D-consistent image synthesis. EG3D [43] introduces an explicit-implicit triplane representation to achieve efficient and high-quality 3D-aware image synthesis.
@@ 3.3 Pose-free 3D Reconstruction
The development of feed-forward models has enabled the reconstruction of 3D scenes from unposed images or videos without the need for per-scene optimization. FlowCam [192] uses a single-view feed-forward NeRF combined with optical flow to estimate poses and fuse multi-view point maps for NeRF-based 3D reconstruction. CoPoNeRF [193] builds and refines 4D correlation maps from image pairs to estimate flow and poses, enabling color and depth rendering.
@@
Additionally, several methods build upon DUSt3R [1] for pose-free 3D reconstruction. DUSt3R itself, as a pioneering feed-forward method, utilizes a transformer-based architecture to regress 3D point maps directly from image pairs. Spann3R [62] augments DUSt3R with a spatial memory network, allowing multiview inputs and improving efficiency by eliminating global alignment. However, Spann3R’s sequential processing introduces error accumulation in reconstruction. Fast3R [59] overcomes this limitation by introducing a global fusion transformer, processing multiple views simultaneously and significantly enhancing reconstruction quality. Conversely, CUT3R [3] refines sequential reconstruction by maintaining and incrementally updating a persistent internal state that encodes scene content.
@@ 3.4 Dynamic 3D Reconstruction
Compared to static scene reconstruction, dynamic scene reconstruction [198] poses significant challenges mainly due to the presence of moving objects, changing viewpoints, and temporal variations in scene geometry. Extending feedforward 3D reconstruction for dynamic scenarios mainly involves robust pose estimation to mitigate moving object interference, together with dynamic area segmentation for updating changing environments.
@@
In addition to pointmap-based dynamic scene reconstruction, several recent methods based on 3D Gaussian Splatting (3DGS) have also been proposed for feed-forward dynamic reconstruction. L4GM [212] proposes the first 4D reconstruction model that produces animated objects from single-view videos using per-frame 3DGS representation. 4D-LRM [213] builds upon a transformer-based large reconstruction model, leveraging data-driven training for dynamic object reconstruction. It draws inspiration from 4D Gaussian Splatting [214] and reconstructs dynamic objects as anisotropic 4D Gaussian clouds. While prior works focus on dynamic object reconstruction, BulletTimer [6] introduces the first feed-forward model for dynamic scene reconstruction. Building on GS-LRM [81], it incorporates a bullet-time embedding into the input frames and aggregates information across all context frames, enabling feed-forward 3D Gaussian Splatting reconstruction at a specific timestamp. In addition, DGS-LRM [215] presents the first feed-forward deformable 3D Gaussian prediction from monocular videos using a transformer-based LRM, while 4DGT [216] extends transformer-based modeling to 4D Gaussian prediction from real-world monocular videos, demonstrating scalability in real-world settings.
@@ 3.6 Image Matching
Recent advances in feed-forward 3D reconstruction have led to significant progress in image matching. One notable example is MASt3R [4], which builds on the DUSt3R [1] to enable efficient and robust image matching in a single forward pass. By augmenting the DUSt3R architecture with a dedicated head for dense local feature extraction, MASt3R introduces a mechanism to improve matching accuracy while maintaining the robustness characteristic of pointmap-based regression. However, MASt3R is fundamentally limited to processing image pairs with poor scalability for large image collections. To address this issue, MASt3R-SfM [230] proposes to leverage the frozen encoder of MASt3R for image retrieval, enabling it to process large and unconstrained image collections with quasi-linear complexity in a scalable way. Importantly, the robustness of MASt3R’s local reconstructions allows the SfM pipeline to dispense with traditional RANSAC-based filtering. Instead, optimization is performed through successive gradient-based refinement in both 3D space (via a matching loss) and 2D image space (via reprojection loss), thus highlighting the potential of feed-forward paradigms to serve as both matching engines and geometric optimizers.
In parallel, to improve generalization for stereo matching, Wen et al. introduce FoundationStereo [231], a large-scale foundation model designed for zero-shot stereo correspondence estimation, pushing the state-of-the-art standard in stereo matching.